OpenStudy (ja1):

-4(2x - 1) = -6(x + 2) - 2

Nnesha (nnesha):

first distribute bracket on left side by -4

OpenStudy (ja1):

\[-8x + 4= -6x - 12 - 2\]

Nnesha (nnesha):

now combine like terms on right side

OpenStudy (ja1):

\[-8x + 4 = -6x - 10\]

Nnesha (nnesha):

-12 and -2 is not 10

OpenStudy (ja1):

woops my bad haha, -14

Nnesha (nnesha):

yes now move -6x on left side

OpenStudy (ja1):

That's where I need help, I'm supposed to divide it correct?

Nnesha (nnesha):

no you have -8x + 4 = - 6x - 14 -6x is subtracting by -14 so what is opposite of subtract ??

OpenStudy (ja1):

So I am supposed to add -6 to both sides?

Nnesha (nnesha):

yes -6x add

OpenStudy (ja1):

\[-2x = -14\]

Nnesha (nnesha):

who kidnapped the "4" :'( where is 4 which is on left side

OpenStudy (ja1):

Wow. I have no idea how I left that out haha: \[-2x + 4 = -14\]

Nnesha (nnesha):

ohh now that looks right now you have to move 4 on right what you should do ???

OpenStudy (ja1):

\[-2x = -18\] \[x = 9\]

OpenStudy (ja1):

Thank so much! It's been a while since I do Algebra and needed a little reminder :)

Nnesha (nnesha):

yep good

Nnesha (nnesha):

my pleasure :) :) :)